# UniErase: Unlearning Token as a Universal Erasure Primitive for Language Models

### v1

@@ 4.1 Unlearning Logical Chain
$\theta\leftarrow\theta+\Delta\theta,\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}\xRightarrow{\mathrm{\pi_{{\theta}}}}y_{\text{idk}}\in\mathcal{D}_{\text{idk}}.$ (6)
This chain enables us to redirect $\mathcal{D}_{f}$ into the unlearning space and leverage model editing to preserve general capabilities. However, the following challenges arise for implementation:
Challenge I: How to teach LLMs to learn and generate [UNL] without overfitting? Since the off-the-shelf soft prompt li2021prefix ; lester2021power ; bhaila2024soft optimization only considers meta tokens as input prefix or suffix without truly introducing a new token. To represent the unlearning space, we need to teach the model the meaning of the unlearning token [UNL]. Besides, editing in Eq 5 requires LLM’s ability to generate the target token (as one step involves maximizing its generation probability). Therefore, we should optimize the model to generate [UNL]. However, the samll $|\mathcal{D}_{f}|$ may lead to overfitting, where the LLM incorrectly holds $q\xRightarrow{\pi_{\theta}}\mathrm{[UNL]}$ for $(q,a)\notin\mathcal{D}_{f}$ (even for $(q,a)\in\mathcal{D}_{r}$) before editing.
Challenge II: How to enhance the robustness of [UNL]? The above derivation is based on model’s continuing generation, but aligned LLMs in practical applications typically append chat templates to input (e.g., user/assistant role, demonstration in Appendix C), which may cause [UNL] fail to induce $\mathcal{D}_{\text{idk}}$. On the other hand, in Eq 5, even if the scale of parameter change is small ($|\Delta\theta|\ll|\theta|$), $\Delta\theta$ may turn Eq 4 not true for the updated model. Thus, for the derivation in Eq 6, we need to ensure that [UNL] remains effective after parameter modifications, successfully serving as a token to induce $y_{\text{idk}}$.
Figure 2: Overview framework of UniErase. We demonstrate current unlearning methods and our UniErase, consisting of unlearning logical chain, unlearning token and unlearning editing.
4.2 Generating Unlearning Token (Proposition 1)
In this section, with Figure 2, we first introduce how to learn the unlearning token [UNL] by introducing a single vector both in the embedding and unembedding matrix of the LLM, and then propose specific techniques to address the two main challenges discussed in the above analysis.
4.2.1 Optimization Target for [UNL] (☞ Challenge I)
Given the embedding matrix of $\pi_{\theta}$ as $E\subset\theta$ and $E\in\mathbb{R}^{n\times d}$, where $n$ is the vocabulary size and $d$ is the embedding dimension, our goal is to add a new token [UNL] and optimize its corresponding embedding vector $e_{u}\in\mathbb{R}^{d}$ to link it with the forgetting space in the LLM via the follow loss:
$\arg\min_{e_{u}}\mathcal{L}_{ul}=\arg\min_{e_{u}}\mathbb{E}_{q\sim\mathcal{D}_{f},a\sim\mathcal{D}_{\text{idk}}}\big{[}-\log p(a\mid q\oplus\mathrm{[UNL]};\pi_{\theta})\big{]},$ (7)
where $p(x\mid y;\pi_{\theta})$ means the conditional probability of $\pi_{\theta}$ generating output $x$ when giving input $y$. Eq 7 ensures that upon encountering [UNL], the model directs the response $a$ (of input $q$) to the forgetting space, with $E\leftarrow E\cup e_{u}\in\mathbb{R}^{(n+1)\times d}$. Prompt Tuning lester2021power shares similar concepts with Eq 7, but to enable the model to generate [UNL], we make some refinements and actually consider:
$\arg\min_{(e_{u},\hat{e}_{u})}\mathbb{E}_{q\sim\mathcal{D}_{f},a\sim\mathcal{D}_{\text{idk}}}\big{[}-\log p(a\oplus\mathrm{[UNL]}\mid q\oplus\mathrm{[UNL]};\pi_{\theta}^{u})\big{]},$ (8)
where $\hat{e}_{u}$ denotes a newly added unembedding vector for [UNL] in the unembedding matrix $U\in\mathbb{R}^{n\times d}$ ($U\leftarrow U\cup\hat{e}_{u}\in\mathbb{R}^{(n+1)\times d}$), and $\pi_{\theta}^{u}$ is the model initializing $e_{u}$ and $\hat{e}_{u}$ randomly. Unlike Eq 7, which only expands $E$, Eq 8 jointly expands both $E$ and $U$. This supervises the LLM to output unlearning expressions $\mathcal{D}_{\text{idk}}$ after [UNL] and equips the model with the ability to generate [UNL] for subsequent edits. However, Eq 8 risks overfitting, as detailed in Chanllenge I in section 4.1.
To address this, we adopt an empirically validated technique: sharing the parameters of [UNL] added to $E$ and $U$ (by setting $e_{u}=\hat{e}_{u}$). This approach has also been employed in the pre-training of some LLMs touvron2023llama . We provide experiments and analysis on solving the overfitting problem in Appendix D.
4.2.2 Robustness Enhancement of [UNL] (☞ Challenge II)
Template Robustness refers to the property where [UNL] can still steer the LLM response toward $\mathcal{D_{\text{ign}}}$ even when additional prefixes/suffixes (such as chat template function $\mathrm{Template}$) are prepended/appended to the original input $q$. Denoting $p_{0}=p(a\mid q\oplus\mathrm{[UNL]};\pi_{\theta})$, we formulate it as:
$\arg\min_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D}_{f}}\big{[}|p_{0}-p(a\oplus\mathrm{[UNL]}|\mathrm{Template}(q\oplus\mathrm{[UNL])};\pi_{\theta})|\big{]}.$ (9)
Parameter Robustness characterizes the feature that after subsequent model edits (as described in Eq 5), [UNL] can still steer the input toward $a\in\mathcal{D_{\text{ign}}}$. Supposing the distribution of parameter perturbation $\epsilon$ is $\mathcal{N}$, the objective can be formalized as follows:
$\arg\min_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D}_{f},\epsilon\sim\mathcal{N}}\big{[}|p_{0}-p(a\oplus\mathrm{[UNL]}\mid q\oplus\mathrm{[UNL]};\pi_{\theta+\epsilon})|\big{]}.$ (10)
Directly solving Eq 9 or 10 is challenging. Therefore, we adopt a reverse approach to enhancing robustness through extra optimization on [UNL] based on Eq 8. Specifically, after initial optimization, we mix $\mathcal{D}_{f}$ and $\mathrm{Template}(\mathcal{D}_{f})$ and optimize $(e_{u},\hat{e}_{u})$ again for Eq 9. Then, since editing only modifies the down-projection matrix $W_{\text{dp}}^{i}$ of the $i$-th MLP layer, for Eq 10, we further optimize [UNL] via adversarial perturbation: $W_{\text{dp}}^{i}\leftarrow W_{\text{dp}}^{i}+\alpha f(W_{\text{dp}}^{i})$, where $f$ is a distribution function.
4.3 Performing Unlearning Edits (Proposition 2)
Upon obtaining [UNL], we then edit the MLP layer of the LLM so that for $\forall(q,a)\in\mathcal{D}_{f}$, the mapping $q\xRightarrow{\pi_{{\theta}}}a$ is modified to $q\xRightarrow{\pi_{{\theta}}}a^{*}\in\mathcal{D}_{\text{idk}}$ for the internalization of unlearning targets. As shown in Figure 2, the $i$-th MLP layer performs the following computation ($\sigma$ and $\gamma$ are activation functions):
@@ 4.3 Performing Unlearning Edits (Proposition 2)
where $h^{i}$, $a^{i}$, and $m^{i}$ represent the hidden state, attention and MLP output in the $i$-th layer, respectively. Unlearning editing only modifies $W_{\text{dp}}^{i}\leftarrow\tilde{W}_{\text{dp}}^{i}$ to ensure $\tilde{W}_{\text{dp}}^{i}k^{*}=v^{*}$ holds, where $k^{*}$ represents the corresponding hidden state of $q$, and $v^{*}$ maximizes the prediction probability of [UNL] as the next token. We provide the details on how to get these new $k^{*}$ and $v^{*}$ for each $(q,a)$ in Appendix E.
In other words, we should then build new mappings from $W_{\text{dp}}^{i}k=v\to W_{\text{dp}}^{i}k^{*}=v^{*}$ to implement unlearning edits described in Eq 5. For each $(q,a)$ in $D_{f}$ and $D_{r}$, we stack their corresponding $k^{*}$, $v^{*}$ vectors into matrices $K_{f}$, $V_{f}$ and $K_{r}$, $V_{r}$, respectively. Following ROME meng2022locating , we build the new mappings by updating $W_{\text{dp}}^{i}\leftarrow W_{\text{dp}}^{i}+\Delta^{*}$ to approximate the overall unlearning standard in Eq 1 via:
$\Delta^{*}=\arg\min_{\Delta}\big{(}\underbrace{\|(W_{\text{dp}}+\Delta)K_{f}-V_{f}\|^{2}}_{\text{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}forget term}}}+\underbrace{\|(W_{\text{dp}}^{i}+\Delta)K_{r}-V_{r}\|^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}retain term}}}\}\big{)}.$ (12)
In Eq 12, for all $(q,a)\in\mathcal{D}_{f}$, the forget term modifies the first token of the response $a$ to [UNL], while the retain term ensures that all $(q,a)\in\mathcal{D}_{r}$ retain their original input-output pairs. In practice, to further maintain model’s ability, we include a sample of general knowledge $\mathcal{D}_{g}$ in $\mathcal{D}_{r}$ ($|\mathcal{D}_{g}|\gg|\mathcal{D}_{f}|$), which other unlearning methods unable to involve due to high computational costs yuan2024closer ; zhang2024negative . Then, to solve Eq 12, we can quickly get its closed-form solution, without optimization process:
@@
In Appendix E, we provide a rigorous mathematical derivation for Eq 32, along with the other unlearning editing methods (derived from 12) that better redirect $D_{f}$ to [UNL] or preserve $D_{r}$.
4.4 Balancing between Unlearning Efficacy and Model Ability
In summary, Sec 4.2 introduces methods to obtain a robust new token as [UNL], clarifying a specific unlearning space. Sec 4.3 then executes unlearning edits to achieve redirecting $\mathcal{D}_{f}$ to [UNL]. Collectively, the reasons that UniErase can balance unlearning and ability lies in:
Effective Unlearning Logic. For unlearning efficay, the unlearning logical chain strictly ensures that unlearning token and unlearning editing work in synergy to achieve the targeted forgetting.
Constrained Parameter Modification. For model ability, unlearning token only adds two new vectors and addresses overfitting, while unlearning editing modifies only several matrices in a few MLP layers while preserving the LLM’s original ability by adding the general knowledge set.
@@ 5 Experiment
In this section, we experimentally validate the universality and effectiveness of our UniErase via multiple unlearning metrics in three scenarios: (I) Batch Unlearning ($\triangleright$ Sec 5.2) refers to enabling an LLM to forget a large forgetting dataset in a single unlearning step, which is a key scenario considered in current domain. (II) Sequential Unlearning ($\triangleright$ Sec 5.3) performs multiple consecutive rounds of batch unlearning, testing whether the LLM collapses under repeated application. (III) Precise Unlearning ($\triangleright$ Sec 5.4) considers extremely small forgetting sets to test the precision of unlearning methods, serving as a practical and meaningful complement to existing scenarios.
@@ 5.1 Overall Settings
Datasets and Models. We consider two widely adopted TOFU and RETURN liu2024learning benchmarks for fictitious and real-world unlearning, respectively. These benchmarks both contain serval forgetting set and corresponding and neighboring retaining set. To test LLMs’ general ability, we further apply MMLU hendrycks2020measuring for fact answering, TriviaQA joshi2017triviaqa for context comprehension, GSM8k cobbe2021training for math reasoning, and Human-Eval chen2021evaluating for coding. Following previous works, we perform unlearning on two scales of Llama-3 touvron2023llama serie model (3B and 8B). Notably, for fictitious unlearning, we use the TOFU-injected  [^1] https://huggingface.co/open-unlearning/tofu_Llama-3.1-8B-Instruct_full version. Details of these datasets and models are demonstrated in Appendix F.
@@
Metrics. We consider multiple metrics across unlearning efficacy and model ability. For unlearning efficacy, in line with prior works maini2024tofu ; zhang2024negative ; yuan2024closer , we employ ROUGE (word-level match), Probability (ground truth likelihood), Truth Ratio (correct-incorrect preference), Token Entropy (generation diversity), Similarity (semantic similarity), and Entailment Score (factual correctness) ferrandez2008te4av . For $\mathcal{D}_{f}$, we use the arithmetic mean of these metrics as the overall Forget Efficacy (FE) yuan2024closer . For model ability, we similarly apply the above metrics to $\mathcal{D}_{r}$, computing their harmonic mean as Retain Efficacy (RE) yuan2024closer . On the other hand, for general abilities, we calculate accuracy (Acc), “I do not know”-like response ratio (Idk), and average response token number (Len) on relevant datasets. Besides, Retain Average (RA) is the mean of all RE and Acc. We provide the details of these metrics in Appendix G.
Baselines. We compare our UniErase with serveral available unlearning methods (detailed in Appendix B): GA liu2022continual , DPO maini2024tofu , NPO zhang2024negative and IDK yuan2024closer , which are the forgetting losses. We further combine them with retaining losses: KL maini2024tofu , GD liu2022continual and ME yuan2024closer , finally forming GA+GD, GA+KL yao2024large , NPO+KL zhang2024negative , ME+GD yuan2024closer , DPO+GD maini2024tofu , DPO+KL, IDK+AP yuan2024closer . We follow each baseline’s parameter settings and provide them in Appendix H. Besides, for the unlearning editing in our UniErase, we apply the AlphaEdit-derived fang2024alphaedit version (detailed in Appendix E).
@@
Table 1: Forget Efficacy (FE), Retain Efficacy (RE) and General Ability of Different Baselines on TOFU benchmark for Batch Unlearning. “Base” means the original LLM before unlearning. In each row of the baselines, we bold the maximum value and underline the second largest one. “Forget” and “Retain” is the most numerous $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ in TOFU, with “Real” as its real fact test set. We provide more experiment results (for the 3B-scale LLM) in Appendix I, showing similar patterns.

### v2

@@ 4.1 Unlearning Logical Chain
$\left(|\Delta\theta|\ll|\theta|,\quad\theta\leftarrow\theta+\Delta\theta\quad\text{s.t.}\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}\right)\quad\land\quad\left(q^{\prime}\xRightarrow{\pi_{{\theta}}}a^{\prime}\right)$ (5)
According to Eq. 5, Udit must demonstrate the ability to alter the subsequent token prediction of target unlearning contents to [UNL] via sparse parameter updates, while maintaining intact knowledge retrieval and response capabilities for non-target contents.
Derivation: Grounded in the aforementioned propositions, we establish the following Unlearning Logical Chain, which directly modifies LLM parameters to accomplish efficient targeted unlearning without compromising the model’s retained knowledge and general capabilities:
$\left(q^{\prime}\xRightarrow{\pi_{{\theta}}}a^{\prime}\right)\quad\land\quad\left(\theta\leftarrow\theta+\Delta\theta\quad\text{s.t.}\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}\xRightarrow{\mathrm{\pi_{{\theta}}}}y_{\text{idk}}\in\mathcal{D}_{\text{idk}}\right)$ (6)
This chain demonstrates the core spirits of UniErase, enabling us to realize unlearning on $\mathcal{D}_{f}$ via directing $a\in\mathcal{D}_{F}$ to $y_{\text{idk}}$, while preserving other untargeted generation $q^{\prime}\to a^{\prime}$.
Figure 2: Paradigm of UniErase and comparison with mainstream FT-based unlearning.
4.2 Unlearning Token
In this section, we present the specific techniques for deriving the unlearning token that fulfill the requirements in the Unlearning Logical Chain. In essence, the special token must satisfy three key criteria: redirecting arbitrary knowledge toward the forgetting space (Eq. 4), maintaining the normal response generation for other knowledge domains (Eq. 4), and being a generatable token (Eq. 5) rather than appearing only at the input end like prefix tuning (Li & Liang, 2021; Bhaila et al., 2024).
4.2.1 Optimization to Attain Unlearning Token
Let $E\in\mathbb{R}^{n\times d}\subset\theta$ and $U\in\mathbb{R}^{n\times d}$ denote the embedding and unembedding matrices of the LLM $\pi_{\theta}$, respectively, where $n$ represents the vocabulary size and $d$ denotes the model dimension. We expand both $E$ and $U$ by incorporating two additional row vectors: $e_{0}\in\mathbb{R}^{d}$ and $u_{0}\in\mathbb{R}^{d}$ ($E\leftarrow E\cup e_{0}$ and $U\leftarrow U\cup u_{0}$), which correspond to the encoding and decoding representations for the unlearning token [UNL], respectively. Subsequently, we optimize the following objective to learn [UNL]:
$\arg\min_{(e_{u},\hat{e}_{u})}\mathbb{E}_{(q,a)\sim\mathcal{D}_{f},a^{\prime}\sim\mathcal{D}_{\text{idk}}}\big[-\log P(a^{\prime}\oplus\mathrm{[UNL]}\mid q\oplus\mathrm{[UNL]};\pi_{\theta}^{u})-\log P(a\mid q;\pi_{\theta}^{u})\big],$ (7)
where $P(x\mid y;\pi_{\theta})$ means the conditional probability of $\pi_{\theta}$ generating output $x$ when giving input $y$. Eq. 7 ensures that upon encountering [UNL], the model directs the original response $a$ (of input $q$) to the forgetting space, yielding $a^{\prime}\in\mathcal{D}_{\text{idk}}$. Concurrently, the LLM is also forced to learn to generate $a^{\prime}\oplus\mathrm{[UNL]}$, satisfying the requirement specified in Eq. 5 that the LLM possesses the capability to generate [UNL]. Furthermore, since we only introduce additional parameters in $E$ and $U$, while keeping all other parameters frozen, the aforementioned [UNL] optimization does not interfere with normal response generation for other knowledge domains.
4.2.2 Robustness Enhancement of Unlearning Token
Parameter Robustness. In the Unlearning Logical Chain, $\Delta$ in Udit may render the previously learned unlearning token ineffective, so we need to enhance its robustness against slight parameter perturbations. While directly incorporating constraints into Eq. 7 to achieve this is challenging, we leverage the fact that Udit’s target parameters are confined to the down projection matrices $W_{dp}$ within the MLP module. Therefore, in the optimization process of Eq. 7, we introduce the following perturbations in the LLM to improve the parameter robustness of the resulting unlearning tokens:
$W_{\text{dp}}^{i}\in\theta\leftarrow W_{\text{dp}}^{i}+\alpha f(W_{\text{dp}}^{i})\cdot W_{\text{dp}}^{i},$ (8)
where parameter $\alpha$ controls the intensity and $f$ is a function mapping $W_{\text{dp}}^{i}$ to a scalar.
4.3 Unlearning Edit
With [UNL] obtained, we propose Udit to bridge it with the knowledge to be unlearned, ensuring the internalization of unlearning targets. Following previous model editing techniques (Meng et al., 2022a; b), we target at the down projection matrices $W_{\text{dp}}^{i}$ in the MLP module. As shown in Figure 2, the $i$-th MLP layer performs the following computation ($\sigma$ and $\gamma$ are activation functions):
@@ 4.3 Unlearning Edit
where $h^{i}$, $a^{i}$, and $m^{i}$ represent the hidden state, attention and MLP output in the $i$-th layer, respectively. Udit exclusively updates $W_{\text{dp}}^{i}\leftarrow\tilde{W}_{\text{dp}}^{i}$ to satisfy the new association $\tilde{W}_{\text{dp}}^{i}k^{*}=v^{*}$, where $k^{*}$ corresponds to the hidden state of targeted knowledge $q\in\mathcal{D}_{f}$, and $v^{*}$ is optimized to maximize the prediction probability of [UNL] as the next token when the input is $q$. In other words, Udit builds new knowledge mappings from the original $W_{\text{dp}}^{i}k=v\to W_{\text{dp}}^{i}k^{*}=v^{*}$ to ensure the next token prediction for $q\in\mathcal{D}_{f}$ is modified to [UNL] (Eq. 5). The detailed procedures for obtaining these $k^{*}$ and $v^{*}$ for each knowledge-answer pair $(q,a)$ are provided in Appendix C.
We construct the unlearning matrices by stacking the corresponding key and value vectors. Specifically, for each $(q,a)\in D_{f}$ to be unlearned, we stack their corresponding unlearning key vectors $k^{*}$ and value vectors $v^{*}$ into matrices ($K_{f}$, $V_{f}$), respectively. Similarly, for each $(q,a)\in D_{r}$ to be retained, compile their normal key and value vectors into matrices ($K_{r}$, $V_{r}$).
Then, we propose the core technique of Udit: by updating $W_{\text{dp}}^{i}\leftarrow W_{\text{dp}}^{i}+\Delta^{*}$, we construct new mappings between $q\in\mathcal{D}_{f}$ and [UNL], while preserving the retrieval of other knowledge ($q\in\mathcal{D}_{r}$) to approximate the unlearning objective described in Eq. 1. The parameter update $\Delta^{*}$ is optimized via:
$\Delta^{*}=\arg\min_{\Delta}\big(\underbrace{\|(W_{\text{dp}}+\Delta)K_{f}-V_{f}\|^{2}}_{\text{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}forget term}}}+\underbrace{\|(W_{\text{dp}}^{i}+\Delta)K_{r}-V_{r}\|^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}retain term}}}\big).$ (10)
In Eq. 10, for all $(q,a)\in\mathcal{D}_{f}$, the forget term modifies the first token of the response $a$ to [UNL], while the retain term ensures that all $(q,a)\in\mathcal{D}_{r}$ retain their original input-output pairs. Through mathematical derivation (provided in Appendix C), we can quickly get its closed-form solution:
@@
Notably, Eq. 11 accommodates a significantly broader $\mathcal{D}_{r}$ than FT-based unlearning methods, thereby preserving a wider range of general knowledge and capabilities. To this end, we include a general knowledge dataset $\mathcal{D}_{g}$ within $\mathcal{D}_{r}$ (with $|\mathcal{D}_{g}|\gg|\mathcal{D}_{f}|$), which remains computationally infeasible for other unlearning approaches (Yuan et al., 2024; Zhang et al., 2024b; Maini et al., 2024).
Null-space Projection Unlearning. Inspired by AlphaEdit (Fang et al., 2024), we further optimize Udit into a null space projection formulation to further reduce the impact of unlearning on general knowledge. Specifically, we obtain the new parameter update $\Delta P$ by right-multiplying with matrix $P$, which projects $\Delta$ onto the null space of $K_{r}$ such that $\Delta PK_{r}=0$. Through straightforward calculation, the retain term in Eq. 10 degenerates to 0 (meaning no influence on the retaining set), thus we only need to optimize the forget term. The new optimization objective can be formulated as:
$\Delta^{*}=\arg\min_{\Delta}\big(\underbrace{\|(W_{\text{dp}}+\Delta P)K_{f}-V_{f}\|^{2}}_{\text{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}forget term}}}+\underbrace{\|\Delta P\|^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}constrain term}}}\big),$ (12)
where the constraint term is incorporated to limit the magnitude of parameter updates. Through a similar derivation as presented in Eq. 11, we arrive at the following closed-form solution:
$\Delta^{*}=(V_{f}-W_{\text{dp}}K_{f})K_{f}^{T}P(K_{f}K_{f}^{T}P+I)^{-1}.$ (13)
The complete mathematical derivation of the closed-form solution presented in Eq. 13, as well as the methodology for computing $P$, is provided in Appendix C.
@@ 5 Experiment
In this section, we experimentally validate and analyze the effectiveness of our balanced and precise UniErase in the following three scenarios: (I) Batch Unlearning ($\triangleright$ Section 5.2) refers to making an LLM forget a large forgetting dataset in a single unlearning step. (II) Sequential Unlearning ($\triangleright$ Section 5.3) performs multiple rounds of unlearning tasks, testing whether unlearning methods cause LLMs to collapse for consecutive scenarios. (III) Precise Unlearning ($\triangleright$ Section 5.4) considers extremely small (single-entry) forgetting sets to test the precision of unlearning methods.
@@ 5.1 Overall Settings
Datasets & Models. We consider two widely adopted TOFU (Maini et al., 2024) and RETURN (Liu et al., 2024) benchmarks for fictitious and real-world knowledge unlearning, respectively. They both contain serval forgetting sets and corresponding and neighboring retaining sets. Highlighting the retention for general abilities, we employ MMLU (Hendrycks et al., 2020) for fact answering, TriviaQA (Joshi et al., 2017) for context comprehension, GSM8k (Cobbe et al., 2021) for math reasoning, and Human-Eval (Chen et al., 2021) for coding. Following previous works (Maini et al., 2024; Yuan et al., 2024), we perform unlearning on the Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Touvron et al., 2023). For fictitious knowledge unlearning, we apply the model versions [^1] https://huggingface.co/open-unlearning/tofu_Llama-3.1-8B-Instruct_full fine-tuned on TOFU. More details of these datasets and models are demonstrated in Appendix D.
@@
Metrics. We consider multiple metrics to comprehensively evaluate performances on unlearning and retention. For unlearning efficacy, in line with prior research (Maini et al., 2024; Zhang et al., 2024b; Yuan et al., 2024), we employ ROUGE (word-level match), Probability (ground truth likelihood), Truth Ratio (correct-incorrect preference), Token Entropy (generation diversity), Similarity (semantic similarity), and Entailment Score (factual correctness) (Ferrández et al., 2008). To obtain an integrated indicator, we calculate the arithmetic mean of these metrics on $\mathcal{D}_{f}$ as the overall Forgetting Efficacy (FE) (Yuan et al., 2024). For neighboring knowledge retention on $\mathcal{D}_{r}$, we similarly apply the above metrics and compute their harmonic mean to be the Retaining Efficacy (RE) (Yuan et al., 2024). Besides, for general abilities, we report the accuracy (Acc), “I do not know”-like response ratio (Idk), and average response token number (Len). Besides, Retaining Average (RA) is the mean of RE and all Accs, as the final metric for retention. We provide the details of these metrics in Appendix E.
Baselines. To demonstrate the effectiveness of our proposed paradigm, we evaluate UniErase against 8 FT-based unlearning baselines. Our comparison includes four primary forgetting losses: GA (Liu et al., 2022; Yao et al., 2024), DPO (Maini et al., 2024), NPO (Zhang et al., 2024b), and IDK (Yuan et al., 2024). These forgetting losses are combined with various retaining loss functions, including KL (Maini et al., 2024), GD (Liu et al., 2022), and ME (Yuan et al., 2024), resulting in the following baseline configurations: GA+GD, GA+KL, NPO+GD, NPO+KL, ME+GD, DPO+GD, DPO+KL, and IDK+AP. The specific parameter settings of all methods are detailed in Appendix F.
@@
Table 1: Batch unlearning performances of different unlearning methods for the TOFU-inject Llama-3.1-8B-Instruct. “Base” means the original LLM before unlearning. In each row, we bold the maximum value and underline the second largest one. “Forget” and “Retain” refer to the $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ datasets in TOFU, while “Real” is the real fact test dataset in TOFU.